OpenStudy (anonymous):

will medal and fan xx Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2

OpenStudy (anonymous):

they both can

OpenStudy (anonymous):

A. x = (x – 6) + 2 B.x + y = y + 2 C.(y + 2) + y = 6 D.x + (y + 2) = 6

OpenStudy (anonymous):

i think it might either be B or D

OpenStudy (igreen):

No, you're wrong.

OpenStudy (igreen):

You're gonna substitute y + 2 for x in the first equation.

OpenStudy (anonymous):

neither one of those are it

OpenStudy (igreen):

That will get you: x + y = 6 (y + 2) + y = 6

OpenStudy (igreen):

Nope.. xD @shortycme

OpenStudy (igreen):

So which option is it? @literally.kaylaa

OpenStudy (anonymous):

OH! I get what you did there! cx Its c Thanks!! @iGreen

OpenStudy (anonymous):

wait im gonna need you to explain it to me @iGreen cuz im confused now

OpenStudy (anonymous):

Lol, yeah^. you basically plug y in to the first equation. but im going to let him explain cause he's better at it :)

OpenStudy (igreen):

The answer is C..because to solve the system of equations by substitution you have to put in x = (y + 2) in the first equation..because the second equation says that x = y + 2. Therefore we can put (y + 2) in the second equation for x: x + y = 6 (y + 2) + y = 6 Get it? @shortycme

OpenStudy (anonymous):

oh i see now

OpenStudy (anonymous):

I have one more question ^.^ but im not going to close this until he gets it . Lol cx

OpenStudy (anonymous):

i got it you can close it now